Solve the Bernoulli Equation "xyy'+y^2=2x"


1
Expert's answer
2022-02-21T16:36:30-0500

Given:

"xyy'+y^2=2x"


We use the substitution

"v=y^2"

Then

"v'=2yy'"

We get the next DE

"\frac{x}{2}v'+v=2x"

"v'+\frac{2}{x}v=4"

"\mu(x)=e^{\int \frac{2}{x}dx}=e^{2 \ln x}=x^2"

"x^2v'+2xv=4x^2"

"(x^2v)'=4x^2"

"x^2v=(4\/3)x^3+C"

"v=(4\/3)x+C\/x^2"

"y=\pm\sqrt{(4/3)x+C/x^2}"